OpenStudy (anonymous):

What is the probability of having 3 children that are all boys?

OpenStudy (anonymous):

assuming that the events "boy or girl" are independent, it would be \(\frac{1}{2}\times \frac{1}{2}\times \frac{1}{2}\)

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

thanks